Please help me solve this word problem about Compounded Account(s):
The Otwell Company puts $18,000 per year into a sinking fund to purchase major equipment.
The fund earns 7.2% interest compounded annually. Find the amount in the fund after 8 years.
(Round your final answer to two decimal places.)
$

Answer by Theo(13342)

if you invest 18000 at the end of each year, then the future value will be equal to 186,011.85.
if you invest 18000 at the beginning of each year, then the future value will be equal to 199,404.70.
the above results were achieved through use of the arachnoid financial calculator found at https://arachnoid.com/finance/
results from the use of that calculator are shown below.
inputs were everything except future value (fv).
output was future value.


i also used the sum of a geometric series formula of Sn = a * (1-r^n)/(1-r).
a is equal 18000.
r is equal to 1.072.
n is equal to 8
the formula becomes Sn = 18000 * (1 - 1.072^8)/(1-1.072) = 186011.85.
this is the same result as 18000 invested at the end of each year.
i would go with 186011.85 as the future value first.
let me know if you have any questions.